OpenStudy (anonymous):

You and some friends are planning a snowboarding holiday, to take place on December 28. The snow depths at three nearby ski resorts on this date in four recent years are shown in the table below, which you have compiled by gathering data from the Internet.

OpenStudy (anonymous):

Thrasher | Xtreme | Gravity Dec. 28, 2000 74 81 54 Dec. 28, 2001 37 34 47 Dec. 28, 2002 56 65 80 Dec. 28, 2003 72 73 57 For example, on December 28, 2002, Mt. Thrasher had 56 inches of snow. Analyze the data and predict which resort will have the best snow for your holiday. A. Gravity Gulch B. Mt. Thrasher C. Xtreme Terror

OpenStudy (anonymous):

@amistre64 @marsss @jim_thompson5910

OpenStudy (anonymous):

HELP!!!

OpenStudy (anonymous):

@nathan917

OpenStudy (nathan917):

For Gravity Gulch, average would be = (74+37+56+72)/4 = 59.75 For Mt. Thrasher, average = (81+34+56+73)/4 = 61 Can you go on from this?

OpenStudy (nathan917):

Just use the same method.

OpenStudy (anonymous):

gravity gulch